OpenStudy (anonymous):

4. If a person weighs 500 N on Earth, what would she weigh on Jupiter, where the acceleration due to gravity is 26 m/s2? Show all work leading to your answer.

OpenStudy (tumblewolf):

First you need to find the mass on earth that would be 490 N = F = ma = (kg)(m/second squared) so F=ma you have F and a so F/a=m so 490/9.8= 50. so her mass is 50 kg. Next we have to find that on Jupiter: F=ma F=(50 kg)(26m/second squared) F= 1300 N

OpenStudy (anonymous):

can you help with another

OpenStudy (tumblewolf):

Sure let's see it

OpenStudy (anonymous):

5. If a freely falling rock were equipped with a speedometer, by how much would its speed readings increase with each second IF it were on a planet where g = 20 m/s2? Show all work leading to your answer.

OpenStudy (tumblewolf):

For the first question, the gravitational acceleration is 9.8m/s^2. So for each second, the speed increases by another 9.8m/sec. You multiply the number of seconds times the acceleration. Cancel the Seconds and you get m/s. 10 sec * 9.8m/s^2 = 98m/s Same for a planet where gravity is 20m/s^2. Speed increases by 20m/s for each second so like in simpler terms V = g*t V = 20*1 = 20 m/s. V = 20*2 = 40 m/s. V = 20*3 = 60 m/s. Increase = 20 m/s with each second.

OpenStudy (anonymous):

thank you so much!

OpenStudy (tumblewolf):

No problem at all! Glad I could help